OpenStudy (anonymous):

How to expand this ? ln [x^2 - x - 2 / (x+4)^2]^1/3

OpenStudy (anonymous):

ln (a^b) = b* ln a 1/3* ln [x^2 - x - 2 / (x+4)^2] Then just expand (x+4)^2 I assume?

OpenStudy (anonymous):

the brackets in the original question are not adequate. the question can not be answered until it is clear also don't forget: \[\log(ab) = \log(a) + \log(b)\]and\[\log(a/b) = \log(a) - \log(b)\]

hartnn (hartnn):

factor x^2-x-2 first then you'll have the form l/3 [ln ab/c^2] which equals, 1/3 [ln a+ln b -2ln c]

OpenStudy (anonymous):

'factor x^2-x-2 first ' The problem is, in this equation the -2 component is followed up by / (x+4)^2 That won't work.

OpenStudy (anonymous):

this is the problem with the improper brackets. they also imply that the ^1/3 is outside of the bracket, therefore not to be put in front to multiply. im pretty sure it's log{[(x^2 - x - 2)/(x+4)^2]^(1/3)} but nobody knows for sure..